# Evidence for Low Universal Equilibrium Black Hole Spin in Luminous Magnetically Arrested Disks

### v1

@@ 2.4 Calculating fluxes
and the specific energy and angular momentum fluxes,
@@
where $dA_{\theta\varphi}=\sqrt{-g}d\theta d\varphi$ is the surface element, and the integrals are over a sphere of radius $r$. When computed at the BH event horizon, these fluxes give the prefactors in Equations (2) and (3), $l_{\rm in}\equiv l(r=r_{\rm H})$ and $e_{\rm in}\equiv e(r=r_{\rm H})$.
$R^{\mu\nu}=\frac{4}{3}E_{\rm rad}u^{\mu}_{\rm rad}u^{\nu}_{\rm rad}+\frac{1}{3}E_{\rm rad}g^{\mu\nu},$ (18)
where $E_{\rm rad}$ is the radiation energy density in the commoving frame and $u^{\mu}_{\rm rad}$ is the radiation 4-velocity. We compute the radiation specific energy and angular momentum fluxes, $e_{\rm rad}(r)$ and $l_{\rm rad}(r)$, in the same way as in Equations (16) and (17), respectively, but where we replace $T^{\mu}_{\nu}$ with $R^{\mu}_{\nu}$. We also define the radiation pressure $p_{\rm rad}=\frac{1}{3}E_{\rm rad}$.
We define the “jet” torque on the BH to be due to the electromagnetic components of the angular momentum and energy fluxes on the BH. We define the “disk” torque on the BH to be due to their hydrodynamic components. We also define the “radiation” torque on the BH due to $e_{\rm rad}(r_{\rm H})$ and $l_{\rm rad}(r_{\rm H})$.
Because numerical floors can add nonphysical mass density and internal energy near the horizon, we apply a magnetization cutoff on the disk torque, following the procedure described in Appendix B of Lowell et al. (2024). We limit the contribution from the disk such that $\sigma=b^{2}/\rho\leq 30$ for modes that use the density floor of $\sigma_{\rm max}=50$. For models with $\sigma_{\rm max}=15$, we choose the disk cutoff of $\sigma=10$. We also apply a $\sigma$-cut to the radiation fluxes, as they are biased by density floors within the BH’s magnetosphere. Note that Ra0.3 and Ra0.94 include this cutoff, but Ra0.4 does not due to data loss.
We can write the total energy extraction efficiency in terms of specific energy flux $e_{\rm in}$ at the BH,
$\eta=\frac{\dot{m}-\dot{E}}{\dot{m}}=1-e_{\rm in}=1-e_{\rm HD}-e_{\rm EM}-e_{\rm rad}.$ (19)
3 Results
Throughout this work, we differentiate simulation families of different disk scale heights by color. Gray points show nonradiative $h/r=0.3$ simulations from Tchekhovskoy et al. (2011), previously analyzed in Lowell et al. (2024). Blue points show $h/r=0.3$ $HR$ models. Purple points show $h/r=0.1$, and green points show $h/r=0.05$ $HR$ models. Either red or orange points show M1 2T-radiation simulations, $h/r=0.03$ $2T$. Finally, black points show the $h/r=0.03$ $1T$, which uses the Noble et al. (2009) cooling scheme.
3.1 Spin-up parameter and equilibrium spin
Figure 2 shows the dimensionless BH spin-up parameter, defined in Equation (20), as a function of BH spin for different disk scale heights. Each point represents a simulation, for which each $s$ value (for a given value of spin) is the sum of the HD, EM, and radiation components:
$s_{\rm MAD}=s_{\rm HD}+s_{\rm EM}+s_{\rm rad}.$ (20)
Radiation torques are only present in simulations with radiation transport, shown in orange. The magnitude of $s$ is an essential part of BH spin-down, as it controls the timescale at which the BH spins down to its equilibrium spin, as we discuss in Section 3.4.
The spin-up parameter for thick MADs is shown in gray, with the gray dashed line representing the semi-analytic model from Lowell et al. (2024) for nonradiative MAD spin-down. When $a\sim-1$, the BH rapidly spins up toward $a=0$ due to a large spin-up parameter, $s$. As $s$ decreases, spin evolution slows until the spin equilibrium is reached at $s=0$. In Figure 2 the vertical lines indicate the equilibrium spin for each case, where the angular momentum and energy fluxes balance, halting further spin evolution. Thick MADs differ from NT disks, particularly at high spin ($a\sim 1$), where $s<0$ indicates BH spin-down. As $a$ decreases, $s$ increases, eventually vanishing, and thus reaching equilibrium at $a_{\rm eq}\approx 0.07$.
In contrast, thinner disks ($h/r=0.1$, shown in purple) exhibit a spin-up parameter with a factor of $\sim 2$ lower magnitude at high spin. This reduction in $s$ results in a higher equilibrium spin, $a_{\rm eq}\simeq 0.29$, where $s(a_{\rm eq})=0$, marked by the purple vertical line. While this is higher than the nonradiative disk value, it remains significantly lower than the NT disk prediction ($a_{\rm eq}\simeq 1$). The purple dashed line represents our new semi-analytic model for spin-down in thin MADs (see Section 3.4).
Orange and green points in Figure 2 show that the magnitude of $s$ remains largely unchanged for our thinner disks ($h/r=0.05$ and $h/r=0.03$) compared to $h/r=0.1$. We find a slightly higher equilibrium spin for our thinnest ($h/r=0.03$) disks, with $a_{\rm eq}\simeq 0.31$. We do not have enough sampling of $a$ to compute the equilibrium spin for $h/r=0.05$, however, we can use the similar trends (i.e. the value of the slope, $ds/da$) found for simulations computed with $h/r=0.03$ and $h/r=0.1$ to extrapolate the value of equilibrium spin for $h/r=0.05$: we find $a_{\rm eq}\simeq 0.3$. This value appears consistent with the comparable disk scale heights, falling between the curves for the thinner ($h/r=0.03$) and thicker ($h/r=0.1$) disks. The values of equilibrium spin for our thinest disks remain considerably lower than what is expected from razor thin NT disks.
Figure 2: BH spin-down in thin MADs is less efficient than in thick MADs, resulting in a higher – but still much less than unity – equilibrium spin, $a_{\rm eq}$, as revealed by the plot of spin-up parameter, $s$, vs. BH spin, $a$. The Novikov-Thorne (NT) disk spin-up curve is shown by the dash-dot blue line. The radiatively-inefficient MADs with $\Gamma=4/3$ are shown by the gray points, their semi-analytic model from Lowell et al. (2024) is shown as the gray dashed curve. The $h/r=0.1$ curve is shown in purple, the $h/r=0.05$ simulation is shown in green, and the $h/r=0.03$ curve is shown in orange for 2T and black for 1T. We also plot higher-resolution radiatively-inefficient MADs with $\Gamma=13/9$ with brown triangles. Equilibrium spins, where each spin-up curve vanishes, are shown with vertical bars of the same color. The values of $a_{\rm eq}$ for all thin disks, $h/r\lesssim 0.1,$ all cluster at $a_{\rm eq}^{\rm thin}\sim 0.3$, which is much smaller than $a_{\rm eq}^{\rm NT}\approx 1$ and much larger than $a_{\rm eq}^{\rm thick}\approx 0.07$.
In Appendix A, we present a resolution study for the thin disks ($h/r=0.1$ and $h/r=0.05$). For $h/r=0.1$, we find that higher resolution does not affect the value of $s$, suggesting they yield the same $a_{\rm eq}$. This implies that in our low resolution $h/r=0.1$ simulations (models H1a#), the values of $s$ are converged. However, for $h/r=0.05$, higher resolution runs lead to significantly smaller values of $s$ and result in smaller $a_{\rm eq}$ values – reduced from $0.4$ to $0.3$. We conclude that a vertical resolution of $\tilde{N}_{\theta}\lesssim 7$, where $\tilde{N}_{\theta}$ is the number of cells per scale height, is insufficient to accurately determine the equilibrium spin. High resolution in the disk proper (e.g., via SMR) is essential for accurately determining the equilibrium spin for thin MADs. By comparing the torques between low- (under-resolved) and high-resolution (resolved) $h/r=0.05$ simulations, we find that the main differences are in the HD components of both $e_{\rm in}$ and $l_{\rm in}$.
Figure 3: Equilibrium BH spin $a_{\rm eq}$ vs. disk scale height $h/r$ for the four values of scale height studied in this work: 0.03, 0.05, 0.1, and 0.3. Data point colors correspondingly match those in Figure 2. As the disk cools from $h/r=0.3$ to $h/r=0.1$, the equilibrium spin increases. However, cooling the disk even further from $h/r=0.1$ to $h/r=0.05$ or $0.03$ does not change the equilibrium spin significantly. The equilibrium spin appears to converged at quadratically low disk thickness. With the black dashed line we plot our fit, $a_{\rm eq}^{\rm fit}=0.31-2.71(h/r)^{2}$.
Figure 3 plots the values of equilibrium spin, $a_{\rm eq}$, vs. disk thermal scale height, $h/r$. Each point corresponds to the equilibrium spin obtained for the family of simulations for that $h/r$: namely, we find the $a_{\rm eq}(h/r)$ dependence as the root of $s(a)=0$ for that value of $h/r$. We find that when $h/r$ decreases from $0.3$ to $0.1$, the equilibrium spin significantly increases from $a_{\rm eq}=0.07$ to $a_{\rm eq}=0.3$. Surprisingly, however, as the thermal scale height decreases further, down to $h/r=0.05$ and $0.03$, the rise in $a_{\rm eq}$ slows down and appears to saturate just above $a=0.3$. To understand this puzzling feature of MAD spin-down, below we investigate: (1) why the equilibrium spin is higher in luminous MADs compared to radiatively-inefficient MADs, and (2) why the equilibrium spin converges as the disk thermal thickness decreases.
3.2 Torques on the BH
Here we unravel how the different BH torque components in our MAD simulations depend on $h/r$. We decompose the total torques into the hydrodynamic and electromagnetic components, and subsequently into the specific fluxes that make up the spin-up parameter.
Figure 4 shows the specific angular momentum flux at the BH horizon vs. BH spin for a range of $h/r$. We show the NT specific angular momentum by the blue dash-dotted line. The gray dashed curves show the specific angular momentum for the thick nonradiative disks of Lowell et al. (2024). We show the angular momentum flux curves for thinner disks, with $h/r=0.03,0.05$, and $0.1$, in red, green, and purple, respectively.
We find that the HD angular momentum flux for thin disks is still considerably lower (by a factor $\sim 2$) than the angular momentum flux of an NT disk. This is consistent with our previous findings that Maxwell torques within the accretion disk lead to smaller than expected HD angular momentum fluxes unto the BH (Lowell et al., 2024). Even if the magnitude is different, the similarity in the spin dependence of $l_{\rm HD}$ to $l_{\rm NT}$ allows us to construct a simple model for this quantity. We show it as the purple dotted dashed line and describe it in Section 3.4. The HD angular momentum in thin disks is slightly larger than in thick disks: $l_{\rm HD}$ can be up to a factor of $\sim 1.4$ times larger in thin disks for negative spins.
Although our thinnest, 1T and 2T $h/r=0.03$ MADs, (black and red open squares, respectively) lie slightly closer to the NT curve—especially when including the angular momentum from radiation in the HD torque (red filled squares)—the difference remains small, reaching at most $25\%$ and averaging below $10\%$.
Contrary to $l_{\rm HD}$, the electromagnetic angular momentum flux, $l_{\rm EM}$, shows a starker difference between thin and thick disks, with at most a factor of $\sim 3$ difference for high BH spin. Even if the magnitudes are remarkably different, the electromagnetic angular momentum flux in thin disks follows a similar trend with $a$ as the thick disk.
For all disk thicknesses, the jets are the main component of angular momentum transport, so understanding $l_{\rm EM}$ is crucial to understanding BH spin-down. Furthermore, as shown above, the hydrodynamic angular momentum flux does not change enough with disk thickness to explain the difference in $s$ for MADs of different thicknesses, as seen in Fig. 2.
Figure 4: Our new analytic model (dashed purple lines) well describes HD and EM components of specific angular momentum supplied by thin MADs across all thicknesses, $0.03\leq h/r\leq 0.1$. Specific angular momentum vs. spin for different disk thermal scale heights: $h/r=0.03$ (2T red and 1T black), $h/r=0.05$ (green), $h/r=0.1$ (purple), and $h/r=0.3$ (gray). [Top panel]: Hydrodynamic angular momentum flux component, $l_{\rm HD}$. MADs exhibit much lower $l_{\rm HD}$ than the NT disk (blue dashed line). Whereas $l_{\rm HD}$ in thick MADs (gray) remains roughly constant, in thinner MADs $l_{\rm HD}$ increases with decreasing $a$. For $h/r=0.1$ (purple), $l_{\rm HD}$ resembles the shape of $l_{\rm NT}$, so we fit it with a (purple dashed) $40\%\,l_{\rm NT}$ curve, capturing the data trend well. For the thinnest disks ($h/r=0.03$), $l_{\rm HD}$ values (open red squares and filled black hexagons) do not deviate significantly from the $h/r=0.1$ results, and lie much lower than $l_{\rm NT}$. Filled red squares show $l_{\rm HD}+l_{\rm rad}$. We note that the $l_{\rm rad}$ measurement for Ra0.4 does not include a $\sigma$-cutoff to account for numerical floors; $\sigma$-cutoffs for Ra0.3 and Ra0.94 indicate that $l_{\rm rad}$ for Ra0.4 would be lower (shown with a red downward arrow). [Bottom panel]: Electromagnetic (EM) angular momentum flux component $l_{\rm EM}$. Angular momentum extraction by jets is strongest for thick MADs (gray) but weaker in cooled MADs ($h/r=0.1$) and remains unchanged for $h/r\leq 0.1$. The semi-analytic EM models for $h/r=0.3$ (gray) and $h/r=0.1$ (purple dashed) are also shown.
Figure 5: The analytic model (dashed purple lines) well describes HD and EM components of specific energy supplied by thin MADs across all thicknesses, $0.03\leq h/r\leq 0.1$. Specific energy flux on the horizon vs. BH spin for varying thermal scale heights: $h/r=0.3$ (gray), $h/r=0.1$ (purple), $h/r=0.05$ (green), and $h/r=0.03$ (2T red and 1T black). [Top panel]: Hydrodynamic (HD) specific energy flux, $e_{\rm HD}$. Cooled MADs with $h/r=0.1$ contribute less HD energy to the BH. The values of $e_{\rm HD}$ for $h/r=0.1$ appear to follow a similar shape vs. $a$ as for the NT disk. We plot $85\%$ the $l_{\rm NT}$ curve with the purple dashed line. The $e_{\rm HD}$ points for $h/r=0.03$ (red and black) follow the same trend as $e_{\rm HD}$ of $h/r=0.1$. Filled red squares show $e_{\rm HD}+e_{\rm rad}$. The $e_{\rm rad}$ measurement for Ra0.4 does not include a $\sigma$-cutoff to account for numerical floors; $\sigma$-cutoffs for Ra0.3 and Ra0.94 indicate that $e_{\rm rad}$ for Ra0.4 would be lower (shown with a red downward arrow). We find little difference in the $e_{\rm HD}$ of thin $h/r=0.1$ and thinnest $h/r\leq 0.05$ MAD. [Bottom panel]: The purple dashed line, well-representing EM specific energy flux thin MADs is simply $1/2$ the thick MAD model (gray dashed line) of Eq. of (28). MADs with spinning BHs extract energy from their BHs (when $e_{\rm EM}<-1$). Thin MADs extract less EM energy than thick MADs. Disks with $h/r\leq 0.05$ extract similar amounts of EM energy as those with $h/r=0.1$.
Figure 5 shows both components of the specific energy flux on the BH horizon, $e_{\rm EM}$ and $e_{\rm HD}$. The color convention for different thermal scale heights is the same as previous Figures. The top panel shows the HD component of energy. The NT disk is shown by the blue dashed line, where it drops as spin increases.
The hydrodynamic energy is lower for thin disks when compared to $e_{\rm NT}$ or $e_{\rm HD}$ for thick MADs. However, as with $l_{\rm HD}$ the difference between thin and thick disks for $e_{\rm HD}$ are at most $50\%$ and are on average closer to $20\%$ for most spins. The $50\%$ difference is due to the fact that, at high spin, unlike the thick MAD which stays roughly constant with spin, the thinner disk follows a similar trend as the NT disk. The fact that $e_{\rm HD}$ for thin disks follows a similar trend to $e_{\rm NT}$ allow us again to model it as a fraction of the NT hydrodynamic energy flux, $e_{\rm HD}=0.85e_{\rm NT}$.
The thinnest 1T and 2T $h/r=0.03$ MADs (black and red open and filled squares) exhibit lower $e_{\rm HD}$ than the $h/r=0.1$ MADs at moderate spin ($a\simeq 0.13$) but reach similar values at $a=0.9375$. However, these differences never exceed $15\%$.
The EM specific energy fluxes also show strong differences between thin and thick disk MAD cases, slightly outshining the differential effect (between thin and thick) of the HD contribution. For high prograde spins, EM energy fluxes are about twice as small in thin MADs as in thick MADs. In contrast, for moderate and retrograde spins, the difference is less significant, with thick MADs showing only about $30-50\%$ stronger fluxes.
3.3 Relationship between spin-down and thermal scale height
In Figure 3, we found that MADs with thinner disks ($h/r=0.1$ or less) have a noticeably higher equilibrium spin ($a\simeq 0.3$) compared to thicker MADs with $h/r=0.3$ ($a\simeq 0.07$).
Moderski & Sikora (1996) showed that torques applied to the BH can be decomposed into torques from the disk and jets. We write this in terms of the spin-up parameter, Equation (5),
$s_{\rm MAD}=l_{\rm HD}-2ae_{\rm HD}-\eta_{\rm EM}\left(\frac{1}{k\Omega_{\rm H}}-2a\right),$ (21)
where we define the EM energy efficiency, $\rm\eta_{\rm EM}=-e_{\rm EM}$, and the EM specific angular momentum flux on the BH, $l_{\rm EM}=\eta_{\rm EM}/(k\Omega_{\rm H})$. Here, $k$ is the ”slippage” parameter, which is the ratio of the field line angular frequency to the angular frequency of the BH horizon [^1] We note that when $a=0$, $\frac{1}{k\Omega_{H}}$ is undefined. However, this does not matter in practice because when $a=0$, all EM energy and angular momentum extraction vanishes., $k=\Omega_{\rm F}/\Omega_{\rm H}$.
Because jets dominate the spin-down process and the disk influences the jet’s opening angle and rotation, we explore the dynamics of jet angular momentum ($k$) and energy ($\rm\eta_{\rm EM}$), and how their interplay with the disk structure determines the equilibrium spin.
3.3.1 Weaker spin-down in thin disks
We start by comparing the jet power and efficiency for different disk thermal thicknesses. In Figure 6 we show EM efficiency $\eta_{\rm EM}$ vs. BH spin with curves for several thermal scale heights. We find that when $h/r$ decreases from $0.3$ to $0.1$, the EM efficiency drops by about a factor of $2$. However, when $h/r$ decreases even further from $0.1$ to $0.03$, the $\eta_{\rm EM}$ decrease stalls, indicating possible convergence. Scepi et al. (2023) and Dhang et al. (2024) also found that $\eta_{\rm EM}$ drops with decreasing thermal scale height, and we find similar values of jet efficiency for $h/r=0.1$. In contrast to Scepi et al. (2023) we find a very different jet efficiency for $h/r=0.03$, with our jets outshining theirs by a factor of $5$, although this might be due to different diagnostics for the jet efficiency.
Figure 6: [Top panel]: Electromagnetic (EM) efficiency vs. BH spin for varying $h/r$. At high spin, the EM efficiency of radiatively-inefficient MADs is roughly twice that of the MADs cooled to a thermal scale height of $h/r=0.1$. For the thinnest disks with $h/r=0.03$, the EM efficiency is roughly the same as the ones with $h/r=0.1$.[Middle panel]: Magnetic flux on the event horizon of the black hole normalized to $\sqrt{\dot{m}}$, also called MADness parameter. We notice than thin MADS have a smaller fluxes at the event horizon. [Bottom panel]: Electromagnetic (EM) power normalized to magnetic flux on the BH horizon vs. BH spin for simulations with varying disk scale height. The gray dot-dashed line is a spline interpolation for $h/r=0.1$ points, included solely to guide the reader’s eye. Although in Figure 6, EM efficiency for $h/r=0.1$ is about 50% that of $h/r=0.3$, the magnetic flux-normalized jet power of all thermal scale heights (or Eddington ratios) are equivalent. Thus, the lower EM power in cooled disks is a result of lower magnetic flux on the BH.
In Figure 6 (middle panel), we show the normalized magnetic flux, $\phi_{\rm BH}=\Phi_{\rm BH}/\sqrt{\dot{m}r_{\rm g}c}$, where

### v2

@@ III.3 Jet efficiency and black hole magnetic flux
To facilitate quantitative comparison to the simulations, let us scale $\phi_{\mathrm{BH}}$ to the representative values, $50$ and $35$, for prograde thick and thin MADs, respectively (see Figure 6b) and the BH rotational rate to a typical value of the BH spin, $a=0.9$ (which corresponds to $\Omega_{\mathrm{H}}r_{\mathrm{g}}\approx 0.3c$ and $f(\Omega_{\mathrm{H}})\approx 1$),
@@
Equations (28) and (29) show that these characteristic values of $\phi_{\mathrm{BH}}$ naturally result in $\eta_{\mathrm{BZ}}$ values of $1$ and $0.5$, for thick and thin MADs at $a=0.9$, respectively, in broad agreement with Fig. 6(a). Thus, we understand rather well how the spin and magnetic flux set the BH spin-down luminosity (at least for prograde MADs). However, the physics that determines the value of $\phi_{\mathrm{BH}}$, as a function of $h/r$, remains poorly understood.
III.4 Angular velocity of the jet field lines
Now we look at the angular momentum extraction due to the EM jets. As we saw in Figure 4, $l_{\mathrm{EM}}$ has a steep dependence on the thermal scale height and changes substantially between $h/r=0.3$ and $h/r=0.1$. Following [63, 54], we can describe the EM specific angular momentum flux on the BH as $l_{\mathrm{EM}}=\eta_{\mathrm{EM}}/(k\Omega_{\mathrm{H}})$ to better understand its behavior. Blandford and Znajek [3] found that the optimal value of $k$ for a jet with monopolar magnetic field geometry is $k\sim 0.5$, as assumed in Eq. (26). Although this value maximizes energy extraction for that specific geometry, it does not necessarily maximize the efficiency for other jet geometries (e.g., parabolic). Deviations from $k\sim 0.5$ impact the efficiency of the angular momentum and energy extraction, with $k=0$ and $k=1$ representing the limiting cases where no energy is extracted. We therefore examine $k$ in our simulations to better understand its dependence on jet geometry, which in turn depends on disk thickness and BH spin.
In Figure 7 we show the parameter $k$ as a function of spin for the different disk thermal thicknesses. We see that the thinner the disk, the larger the value of $k$. A higher $k$, indicating faster rotation of field lines, results in less efficient angular momentum extraction and thus a lower $l_{\mathrm{EM}}$. For a MAD of a given $h/r$, this leads to a higher equilibrium spin.
The rotation velocity of the magnetic field lines, $k$, is tied to the jet structure, making it useful to examine the magnetic field lines connected to the BH in order to understand BH spin-down. The jet structure should in turn depend on the forces acting on the disk and their balance. We examine the jet shape and force balance in Section V, including the link between $k$ and the jet geometry (monopolar vs parabolic) in section V.1. But first, in Section IV, we use the dependencies of $\eta_{\mathrm{EM}}$ and $k$ on $a$ to construct a semi-analytical model of spin evolution for thin MADs.
Figure 7: Thinner disks produce more rapidly rotating jets, as seen in a plot of $k=\Omega_{\mathrm{F}}/\Omega_{\mathrm{H}}$ vs BH spin. The gray points show the values calculated in Lowell et al. [54], and the colored points show cooled disks of various scale heights (see legend). For any value of spin, thick, radiatively inefficient disks have jets with the lowest angular frequency. As the disk cools, the value of $k$ increases and appears to converge as the disk becomes thinner.
IV Semi-analytic model for BH spin-down in thin MADs
Lowell et al. [54] developed a physically motivated semi-analytic model that describes BH spin-down in radiatively inefficient MADs. Using our new suite of luminous MAD simulations with varying $h/r$, we now construct a model for spin-down of thinner MADs.
In panels (a) of both Figures 4 and 5, we find that $l_{\mathrm{HD}}\simeq 0.4l_{\mathrm{NT}}$ and $e_{\mathrm{HD}}\simeq 0.85e_{\mathrm{NT}}$. Thus, we can model the hydrodynamic disk component of thin MADs as follows: $s_{\mathrm{HD,thin}}=l_{\rm HD}-2ae_{\rm HD}=0.4l_{\mathrm{NT}}-0.85e_{\mathrm{NT}}\times 2a$.
The electromagnetic jet component also follows an interesting trend that is straightforward to model. As we showed in Section III.3, $\eta_{\mathrm{EM}}$ for $h/r=0.1$ is approximately $1/2$ of $\eta_{\mathrm{EM}}$ for thick MADs. Thus, we can use the fit from Lowell et al. [54] and model $\eta_{\mathrm{EM}}$ for thin MADs as $\eta_{\mathrm{EM}}^{\mathrm{thin}}=\frac{1}{2}\eta_{\mathrm{EM}}^{\mathrm{thick}}$, where